OpenStudy (anonymous):

find y intercept for 4x+y=-8

OpenStudy (anonymous):

first solve the equation for y

OpenStudy (anonymous):

then use the formula y = mx+b m is the slope b is the y-intercept

OpenStudy (anonymous):

set the x in to zero and solve for y

OpenStudy (anonymous):

4x+y = -8 4*0 + y = -8 0+y = -8 y = -8

OpenStudy (anonymous):

Thank you